OpenStudy (anonymous):

if you have 3 over radical 2, how do you simplify that?

OpenStudy (jdoe0001):

$$ \cfrac{3}{\sqrt{2}} $$ is "simplified" by getting rid of the root in the denominator

OpenStudy (jdoe0001):

$$ \cfrac{n}{\sqrt{m}} \implies \cfrac{n}{\sqrt{m}} \times \cfrac{\sqrt{m}}{\sqrt{m}} \implies \cfrac{n\sqrt{m}}{(\sqrt{m})^2} $$

OpenStudy (jdoe0001):

and \( (\sqrt{m})^2\) is well, "m", so the root is no longer there in the denominator